# TALC: Time-Aligned Captions for Multi-Scene Text-to-Video Generation

### v3

@@ Abstract
Recent advances in diffusion-based generative modeling have led to the development of text-to-video (T2V) models that can generate high-quality videos conditioned on a text prompt. Most of these T2V models often produce single-scene video clips that depict an entity performing a particular action (e.g., ‘a red panda climbing a tree’). However, it is pertinent to generate multi-scene videos since they are ubiquitous in the real-world (e.g., ‘a red panda climbing a tree’ followed by ‘the red panda sleeps on the top of the tree’). To generate multi-scene videos from a pretrained T2V model, we introduce Time-Aligned Captions (TALC) framework. Specifically, we enhance the text-conditioning mechanism in the T2V architecture to recognize the temporal alignment between the video scenes and scene descriptions. As a result, we show that the pretrained T2V model can generate multi-scene videos that adhere to the multi-scene text descriptions and be visually consistent (e.g., w.r.t entity and background). Our TALC-finetuned model outperforms the baseline methods on multi-scene video-text data by 15.5 points on aggregated score, averaging visual consistency and text adherence using human evaluation. The project website is https://talc-mst2v.github.io/
Figure 1: Multi-scene video generation methods. (a) Merging scene descriptions; (b) Generating a video per scene independently and concatenating the resulting videos.; (c) TALC: conditioning the generated video on scene-1 description for the first half of the video frames and on scene-2 description for the later video frames.
@@ 1 Introduction
The ability to generate videos that simulate the physical world has been a long-standing goal of artificial intelligence [1, 2, 3, 4]. In this regard, text-to-video (T2V) models have seen rapid advancements by pretraining on internet-scale datasets of images, videos, and texts [5, 6, 7, 8, 9]. Specifically, latest T2V models such as ModelScope [10], Lumiere [11], VideoCrafter [12, 13], and OpenSora [14], train a diffusion model [15] with single-scene video-text datasets [16, 17] that are widely available and easy to acquire. As a result, these models can only generate single scene videos that depicts a single action or event [^1] While high-performance T2V models such as Sora [4] might be able to generate multi-scene videos, we point out that they are closed-source models trained with massive compute resources and lack sufficient details on the model design, training protocol, and datasets.. However, real-world scenarios (e.g., movie and ads creation) often require the generation of multi-scene video for text descriptions (e.g., Scene1: ‘A koala is napping on a tree.’ Scene2: ‘The koala eats leaves on the tree.’). In such cases, the generated video should accurately depict the events mentioned in the multi-scene script while maintaining visual consistency (e.g., entity and background consistency).
@@
Multi-scene T2V generation differs from long video synthesis where the goal is to either interpolate (few frames to many frames) [7] or create continuing patterns of the single event in the generated video [8]. Prior works [18, 19] use a transformer [20, 21] to generate video frames for a given scene autoregressively. However, these models straggle to generate multiple scenes reliably as the context length increases with the history of text descriptions and visual tokens [22] of the previous generated videos, e.g., generating Scene 4 conditioned on Scene1, 2, 3 videos and descriptions. Some works [23] utilize a latent diffusion model [24] to generate video frames autoregressively by conditioning on the entire history of generated videos and scene descriptions. However, this approach is slow due to repeated sampling, generates only one frame per scene description, and is shown to work with only limited cartoon characters [25, 26] and not with a wide range of visual concepts in the real-world. Other works [27, 28] generate multi-scene videos by utilizing knowledge of the entity, background, and their movements from large language models [29]. However, these videos are generated independently for each scene before being merged. Moreover, these methods do not offer a way to learn from real-world multi-scene video-text data.
@@
To address these challenges, we propose TALC (Time-ALigned Captions), a simple and effective framework to generate consistent multi-scene videos by effectively leveraging the capabilities of pretrained T2V diffusion models. As shown in Figure 1(c), TALC conditions the visual representations of earlier video frames on the embeddings of the earlier scene description, and likewise, it conditions the representations of later video frames on the embeddings of the later scene description in the temporal dimension. The temporal modules in the T2V diffusion architecture allows information sharing between video frames (the first half and the second half) to maintain visual consistency. Thus, TALC enhances the scene-level text-video alignment while providing all the scene descriptions to the diffusion model at once. Further, our TALC framework can enhance the multi-scene text-to-video generation capabilities with real-world multi-scene data (§3.2). We provide a qualitative example for comparing different multi-scene video generation methods in Appendix §C.
In our experiments, we assessed the visual consistency (background and entity consistency) and multi-scene script adherence of the generated videos from Modelscope [10] and Lumiere [11]. In addition, to our TALC method, we evaluated concatenation approaches, either merging descriptions of different scenes or generating a video per scene independently and concating the output videos (Figure 1(a,b)). Through our automatic and human evaluation, we found that merging scene descriptions leads to high visual consistency but poor text adherence. On the other hand, merging videos independently achieves high text adherence but visual consistency is compromised. Our TALC model strikes an effective balance between visual consistency and text adherence, outperforming the baseline methods by $11.1$ points on the overall score, averaging visual consistency and text adherence scores, as determined by human evaluation. Furthermore, we construct a multi-scene text-video dataset from real-world videos and fine-tune the T2V generative model using TALC. On our human evaluation, the generated videos from the TALC-finetuned model exhibit higher text adherence than the base model in multi-scene scenarios. Specifically, it outperforms the baseline methods by $15.5$ points on the overall score.
@@ 2 Preliminaries
In this work, we focus on generating multi-scene videos from text descriptions using a diffusion-based Text-to-Video (T2V) generative model. The initial step is to equip the generative model with the knowledge of a wide range of visual concepts and actions. This is achieved during the pretraining stage (§2.1). Subsequently, we aim to utilize the pretrained model for multi-scene text-to-video generation task, which we formalize in (§2.3).
@@ 2.1 Diffusion Models for Text-to-Video Generation
Diffusion models [15, 30] $p_{\theta}(x)$ are a class of generative models that learn data distribution $p_{data}(x)$. Their class-conditional versions can be trained to learn conditional distributions $p_{data}(x|y)$ where $y$ is the conditioning variable, which could take various forms such as labels from a dataset or text description accompanying a video [31].
We assume a dataset $\mathcal{S}\subset\mathcal{V}\times\mathcal{T}$ consisting of pairs of $(V_{j},T_{j})$ where $V_{j}\in\mathbb{R}^{L\times 3\times H\times W}$ is a raw video consisting of $3$ RGB channels, $L$ frames, $H$ height, $W$ width, and $T_{j}$ is a text caption. The T2V model is trained to learn the conditional distribution of the videos conditioned on the text $p_{\mathcal{S}}(V_{j}|T_{j})$. In this work, we consider diffusion-based generative models that learn this distribution via iterative denoising of the video $z_{j}\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$. Here, $z_{j}$ can either represent the video in the raw pixel space $V_{j}$ [11] or it can represent the latent representation of the video $z_{j}=\mathcal{E}(V_{j})$ for the latent diffusion models [24] where $\mathcal{E}$ is an encoder network such as VAE [32]. Given $z_{j}$, diffused variable $z_{\tau,j}=\alpha_{\tau}z_{j}+\beta_{\tau}\epsilon$ are constructed where $\epsilon\sim\mathcal{N}(0,I)$ and $\alpha_{\tau}$ and $\beta_{\tau}$ are sampled from the noise scheduler $p_{\tau}$ [33], which defines the noise levels the model is trained on. Finally, a denoiser network $f_{\theta}$ [34, 35], which takes the diffused variable $z_{\tau}$ and embeddings of the text caption as input, is trained to predict the target vector $y$ where $y$ can be the original noise $\epsilon$, which minimizes the denoising score matching objective [15]:
@@
where $h_{j}=\mathcal{H}(T_{j})\in\mathbb{R}^{d}$ is the embedding of the text caption $T_{j}$, $\mathcal{H}$ is the text embedding model [36] and $d$ is the dimension size.
@@ 2.2 Text Conditioning Mechanism
For effective textual controllability of video generation, the denoiser networks is equipped with a cross-attention mechanism [10, 7]. Specifically, it conditions the visual content $z_{\tau}\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$ on the text. To do so, we first repeat the text embeddings of the text caption $r_{j}=R(h_{j})\in\mathbb{R}^{L\times d}$ where $R$ is a function that repeats the input text embedding $h_{j}$ for $L$ times in the temporal dimension. Intuitively, the repeat operation represents that the $L$ frames of the video $z_{j}$ are semantically aligned with the textual description $T_{j}$ or its text embedding $r_{j}$. In §3, we will manipulate this operation to make the model architecture aware of the video-text alignment in the multi-scene scenario.
These repeated text embeddings $r_{j}$ are inputs to the spatial attention block as the key and value in the multi-head attention block. The cross-attention enables the intermediate visual features to capture the semantic information that facilitates an alignment between the text and vision embeddings. Formally,
@@
where $CA_{f_{\theta}}$ is the cross attention mechanism with $Q,K,V$ as the query, key, and value, respectively, in the spatial blocks of the denoiser network, $z^{\prime}_{\tau,j}$ is the intermediate representation that is informed with the visual and textual content of the data. In addition to the spatial blocks, the denoiser network also consists of temporal blocks that aggregate features across video frames, which are useful for maintaining visual consistency in the generated video.
@@ 2.3 Multi-Scene Text-to-Video Generation
We aim to generate multi-scene video $X=\{x_{1},x_{2},\ldots,x_{n}\}$ from multi-scene descriptions $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ where $n$ are the number of scenes and each sentence $y_{j}$ is a description for scene $j$. Index $j$ also defines the temporal order of events in the multi-scene script, i.e., if $k>j$, we want the events described in scene $j$ to be depicted earlier than the events described in scene $k$. Further, we want that each video part $x_{j}$, within the entire generated video $X$, to have high video-text semantic alignment with its corresponding scene description $y_{j}$.
Figure 2: Time-Aligned Captions (TALC) Framework. During the generation process of the video, the initial half of the video frames are conditioned on the embeddings of the description of scene 1, while the subsequent video frames are conditioned on the embeddings of the description of scene 2.
@@ 3 TALC: Time-Aligned Captions for Multi-Scene T2V Generation
In this section, we present our novel method TALC, describing its architecture and a novel approach for constructing synthetic multi-scene video-text data to train TALC.
3.1 TALC Architecture
Our starting point is that each specific scene $x_{j}$ in a multi-scene video $X\in\mathbb{R}^{L\times 3\times H\times W}$ should reflect the entities and events described in the conditioning scene description $y_{j}$. To achieve this, TALC adapts T2V diffusion architecture $f_{\theta}$ to multi-scene generation by ensuring that the representation for video part $x_{j}$ attends and aggregates language features specifically from description $y_{j}$. Concretely, TALC allocates $L/n$ frames within the entire video to each $x_{j}\in X$. Let $z_{X}=[z_{x_{1}};z_{x_{2}};\ldots;z_{x_{n}}]\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$ be the representation for the entire video $X$, and $z_{x_{j}}\in\mathbb{R}^{(L/n)\times C\times H^{\prime}\times W^{\prime}}$ for the different scenes, which are concatenated in the temporal dimension. In addition, $r_{Y}=\{r_{y_{1}},\ldots,r_{y_{n}}\}$ be the sequence of text embeddings for the multi-scene description $Y$ and $y_{j}$ be an individual scene description. Under TALC, Eq. 2 is changed to:
@@ 3.1 TALC Architecture
with, $\tau$ the timestamp in the diffusion modeling setup, which is applied during training as well as inference. We illustrate the framework in Figure 2.
@@
While TALC aims to equip the generative model with the ability to follow all the events in the multi-scene descriptions, the visual consistency is ensured by the temporal modules (attentions and convolution blocks) in the existing denoiser network. By design, our approach can be applied to the pretrained T2V model during inference.
3.2 TALC Training
We note that the text adherence capabilities of the pretrained T2V generative model are limited. This is due to the lack of multi-scene video-text data during pretraining, since these datasets are not widely available and are hard to curate for model training. Prior work, such as ActivityNet [37], curated human captions for specific video scenes depicting useful actions in long videos. However, these scenes are either overlapping or have a large temporal gap between them, which are less suitable for natural and smooth transitions required in multi-scene videos.

### v4

@@ Abstract
Most of these text-to-video (T2V) generative models often produce single-scene video clips that depict an entity performing a particular action (e.g., ‘a red panda climbing a tree’). However, it is pertinent to generate multi-scene videos since they are ubiquitous in the real-world (e.g., ‘a red panda climbing a tree’ followed by ‘the red panda sleeps on the top of the tree’). To generate multi-scene videos from the pretrained T2V model, we introduce a simple and effective Time-Aligned Captions (TALC) framework. Specifically, we enhance the text-conditioning mechanism in the T2V architecture to recognize the temporal alignment between the video scenes and scene descriptions. For instance, we condition the visual features of the earlier and later scenes of the generated video with the representations of the first scene description (e.g., ‘a red panda climbing a tree’) and second scene description (e.g., ‘the red panda sleeps on the top of the tree’), respectively. As a result, we show that the T2V model can generate multi-scene videos that adhere to the multi-scene text descriptions and be visually consistent (e.g., entity and background). Further, we finetune the pretrained T2V model with multi-scene video-text data using the TALC framework. We show that the TALC-finetuned model outperforms the baseline by achieving a relative gain of $29\%$ in the overall score, which averages visual consistency and text adherence using human evaluation.
Figure 1: Examples of multi-scene video generation using the TALC framework. Our proposed method, TALC, enhances scene-level alignment between text and video, facilitating the generation of videos with seamless transitions between textual descriptions while preserving visual consistency. The first two rows are generated using Lumiere [6] and the last two rows use ModelScope [49].
@@ 1 Introduction
The ability to generate videos that simulate the physical world has been a long-standing goal of artificial intelligence  [1, 47, 10, 36]. In this regard, text-to-video (T2V) models have seen rapid advancements by pretraining on internet-scale datasets of images, videos, and texts [9, 6]. Previous work [20, 18, 28, 49, 8, 7] primarily focus on training conditional denoising diffusion probabilistic models [22] on paired video-text data [4, 54]. After training, these models allow for video generation by sampling from the trained diffusion model, conditioned on a text prompt. However, most of the open-models such as ModelScope[49] VideoCrafter [13, 14], OpenSora [57] are trained with single-scene video-text dataset [4, 50], which is widely available and easy to acquire. However, real-world scenarios often require the generation of multi-scene videos from multi-scene descriptions (e.g., Scene1: ‘A koala is napping on a tree.’ Scene2: ‘The koala eats leaves on the tree.’). In such cases, the generated video should accurately depict the events in their temporal order (e.g., Scene2 follows Scene1) while maintaining visual consistency, meaning that backgrounds and entities should remain consistent across scenes. While high-performance T2V models such as Sora [36] might be able to generate multi-scene videos, we point out that they are closed-source models trained with massive compute resources and lack sufficient details on the model design, training protocol, and datasets. In this work, we present a complementary approach and tackle the challenge of effectively leveraging the capabilities of base T2V models for multi-scene video generation.
@@
The multi-scene T2V generation differs from long video synthesis where the goal is to either interpolate (few frames to many frames) [18] or create continuing patterns of the single event in the generated video [8]. Prior work like Phenaki [45, 28] use transformers [44, 3] to generate video frames for a given scene autoregressively. However, it is hard for their model to generate multiple scenes reliably as the context length increases with the history of text descriptions and visual tokens [56] of the previous generated videos (e.g., generating Scene 4 conditioned on the Scene1, 2, 3 videos and descriptions). Other works [40] utilize a latent diffusion model [41] to generate video frames autoregressively by conditioning on the entire history of generated videos and scene descriptions. However, the approach is (a) slow due to repeated sampling, (b) generates only one frame per scene description, and (c) shown to work with only limited cartoon characters [29, 55] instead of wide range of visual concepts in the real-world. In this work, our goal is to generate multi-scene videos in the end-to-end manner using a diffusion T2V generative model. Prior work like VideoDirectorGPT [30, 31] generates multi-scene videos by utilizing knowledge of the entity, background, and their movements from large language models [2]. However, these videos are generated independently for each scene before being merged.
@@
To remedy these challenges, we propose TALC (Time-ALigned Captions), a simple and effective framework to generate consistent and faithful multi-scene videos. In particular, our approach conditions the T2V generative model with the knowledge of the temporal alignment between the parts of the multi-scene video and multi-scene descriptions (Figure 3). Specifically, TALC conditions the visual representations of earlier video frames on the embeddings of the earlier scene description, and likewise, it conditions the representations of later video frames on the embeddings of the later scene description in the temporal dimension. Additionally, the temporal modules in the T2V diffusion architecture allows information sharing between video frames (the first half and the second half) to maintain visual consistency. Therefore, TALC enhances the scene-level text-video alignment while providing the scene descriptions to the diffusion model all at once (Figure 1).
Prior methods like FreeNoise [38] propose a motion injection strategy to address these challenges. However, this method is quite sophisticated and difficult to control due to diverse hyperparameters, such as time-specific motion injection, prompt interpolation, and injection in specific cross-attention layers. In contrast, our approach eliminates these complexities and introduces a straightforward mechanism that significantly boosts performance. Unlike previous work, we also demonstrate that finetuning with TALC using real-world multi-scene data can enhance generation capabilities. Specifically, we propose a pipeline to curate multi-scene video data and subsequently finetune the T2V model on this multi-scene data using TALC3.2).
Figure 2: Summary of the results. We compare several baselines and our TALC framework with ModelScope [49] for generating multi-scene videos on the unseen prompts. Specifically, we study the overall score which averages the visual consistency (object and background) and text adherence (video-text alignment). We observe that using TALC with the base model i.e., training-free achieves relative gains of $21\%$ in comparison to FreeNoise [38]. In addition, we show that multi-scene finetuning with TALC (TALC-FT) allows larger gains, achieving relative gains upto $29\%$.
In our experiments, we assess the visual consistency (background and entity consistency) and multi-scene script adherence of the generated videos from T2V generative models. Through our human evaluation, we find that TALC strikes an effective balance between visual consistency and text adherence, and outperforms the FreeNoise by achieving relative gains of $21\%$ points on the overall score. This score represents the average of visual consistency and text adherence scores (Figure 2). Furthermore, we construct a multi-scene text-video dataset from real-world videos and fine-tune the T2V generative model using TALC. We show that finetuning with TALC outperforms FreeNoise by achieving relative gains of $29\%$ on the overall score (Figure 2). Further, we perform automatic evaluation for scalable judgements (§5.4) and provide qualitative examples to highlight the benefits of our approach (§5.5). We present the related work in Appendix §A.
@@ 2.1 Diffusion Models for Text-to-Video Generation
Diffusion models [22, 33] $p_{\theta}(x)$ are a class of generative models that learn data distribution $p_{data}(x)$. Due to their flexible design, we can train their class-conditional versions to learn class-conditional data distributions $p_{data}(x|y)$ where $y$ is the conditioning variable. We assume a dataset $\mathcal{S}\subset\mathcal{V}\times\mathcal{T}$ consisting of pairs of $(V_{j},T_{j})$ where $V_{j}\in\mathbb{R}^{L\times 3\times H\times W}$ is a raw video consisting of $3$ RGB channels, $L$ frames, $H$ height, $W$ width, and $T_{j}$ is a text caption. We use $\mathcal{V}$ and $\mathcal{T}$ to denote the domain of videos and text, respectively. The aim of T2V generative modeling is to learn the conditional distribution of the videos conditioned on the text $p_{\mathcal{S}}(V_{j}|T_{j})$. In this work, we consider diffusion-based generative models that learn the data distribution via iterative denoising of the input video $z_{j}\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$. Here, $z_{j}$ can either represent the input video in the raw pixel space $V_{j}$ [6] or it can represent the latent representation of the video $z_{j}=\mathcal{E}(V_{j})$ for the latent diffusion models [41] where $\mathcal{E}$ is an encoder network [27].
Given $z_{j}$, diffused variable $z_{\tau,j}=\alpha_{\tau}z_{j}+\beta_{\tau}\epsilon$ are constructed where $\epsilon\sim\mathcal{N}(0,I)$ where $\alpha_{\tau}$ and $\beta_{\tau}$ are sampled from the noise scheduler $p_{\tau}$ [16]. Finally, we train a denoiser network $f_{\theta}$ [42, 37] that inputs the diffused variable $z_{\tau}$ and embeddings of the text caption to predict the target vector $y$ where $y$ can be the original noise $\epsilon$, which minimizes the denoising objective [22]:
@@
where $h_{j}=\mathcal{H}(T_{j})\in\mathbb{R}^{d}$ is the embedding of the text caption $T_{j}$ where $\mathcal{H}$ is the text embedding model [39] and $d$ is the dimension size.
@@ 2.2 Text Conditioning Mechanism
To ensure the effective textual controllability of video generation, the structure of the denoiser networks is equipped with a cross-attention mechanism [49, 18]. Specifically, it conditions the visual content $z_{\tau}\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$ on the text. To do so, we first repeat the text embeddings of the text caption $r_{j}=R(h_{j})\in\mathbb{R}^{L\times d}$ where $R$ is a function that repeats the input text embedding $h_{j}$ for $L$ times in the temporal dimension. Intuitively, the repeat operation represents that the $L$ frames of the video $z_{j}$ are semantically aligned with the textual description $T_{j}$ or its text embedding $r_{j}$.
These repeated text embeddings $r_{j}$ are inputs to the spatial attention block as the key and value in the multi-head attention block. The cross-attention enables the intermediate visual features to capture the semantic information that facilitates an alignment between the language and vision embeddings.
@@
where $CA_{f_{\theta}}$ is the cross attention mechanism with $Q,K,V$ as the query, key, and value, respectively, in the spatial blocks of the denoiser network. Additionally, $z^{\prime}_{\tau,j}$ is the intermediate representation that is informed with the visual and textual content of the data. In addition to the spatial blocks, the denoiser network also consists temporal blocks that aggregate features across video frames which are useful for maintaining visual consistency in the video.
@@ 2.3 Multi-Scene Text-to-Video Generation
In this work, we aim to generate multi-scene video $X=\{x_{1},x_{2},\ldots,x_{n}\}$ from multi-scene descriptions $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ where $n$ are the number of sentences and each sentence $y_{j}$ is a scene description for scene $j$. Additionally, the index $j$ also defines the temporal order of events in the multi-scene script i.e., we want the events described in the scene $j$ to be depicted earlier than the events described in the scene $k$ where $k>j$. Further, we want the parts of the entire generated video $X$, given by $x_{j}$, to have high video-text semantic alignment with the corresponding scene description $y_{j}$. In addition, we expect the appearance of the objects to remain consistent throughout the video unless a change is specified in the description.
@@ 3 TALC: Time-Aligned Captions for Multi-Scene T2V Generation
3.1 Approach
Most of the existing T2V generative models [49, 13, 6] are trained with large-scale short video-text datasets (10 seconds - 30 seconds) such as WebVid-10M [4]. Here, each instance of the dataset consists of a video and a human-written video description. These videos either lack the depiction of multiple events, or the video descriptions do not cover the broad set of events in the video, instead focusing on the major event shown. As a result, the pretrained T2V generative models only synthesize single video scenes depicting individual events.
We introduce TALC, a novel and effective framework to generate multi-scene videos from diffusion T2V generative models based on the scene descriptions. Our approach focuses on the role of text conditioning mechanism that is widely used in the modern T2V generative models (§2.2). Specifically, we take inspiration from the fact that the parts of the generated video $x_{j}$ should depict the events described in the scene description $y_{j}$. To achieve this, we ensure that the representations for the part of the generated video aggregates language features from the scene description $y_{j}$.
Consider that we want to generate a multi-scene video $X\in\mathbb{R}^{L\times 3\times H\times W}$ from the scene descriptions $y_{j}\in Y$, using a T2V generative model $f_{\theta}$. Furthermore, we assume that individual video segments $x_{j}$ are allocated $L/n$ frames within the entire video $X$. Let $z_{X}=[z_{x_{1}};z_{x_{2}};\ldots;z_{x_{n}}]\in\mathbb{R}^{L\times C\times H^{\prime}\times W^{\prime}}$ represent the representation for the entire video $X$, and $z_{x_{j}}\in\mathbb{R}^{(L/n)\times C\times H^{\prime}\times W^{\prime}}$ for the $j^{th}$ part of the video that are concatenated in the temporal dimension. In addition, consider $r_{Y}=\{r_{y_{1}},\ldots,r_{y_{n}}\}$ be the set of text embeddings for the multi-scene description $Y$ and $y_{j}$ be an individual scene description. In the TALC framework, the Eq. 2 is changed to:
@@ 3.1 Approach
Here, $\tau$ represents the timestamp in the diffusion modeling setup, which is applied during training as well as inference. We illustrate the framework in Figure 3. While TALC aims to equip the generative model with the ability to depict all the events in the multi-scene descriptions, the visual consistency is ensured by the temporal modules (attentions and convolution blocks) in the denoiser network. By design, our approach can be applied to the base T2V model without any further training.
Figure 3: The architecture of Time-Aligned Captions (TALC). During the generation process of the video, the initial half of the video frames are conditioned on the embeddings of the description of scene 1 ($r_{y_{1}}$), while the subsequent video frames are conditioned on the embeddings of the description of scene 2 ($r_{y_{2}}$).
3.2 Multi-Scene Video-Text Data Generation